# R-AVST: Empowering Video-LLMs with Fine-Grained Spatio-Temporal Reasoning in Complex Audio-Visual Scenarios

### v1

@@ Introduction
The rapid advancement of multimodal large language models (MLLMs) recently has demonstrated their effectiveness in video understanding by integrating information from various modalities (Wang et al. 2024b; Zhang et al. 2024; Tang et al. 2025a; Geng et al. 2025). However, this rapid progress raises a question: Do current models and datasets adequately account for the compositional audio-visual nature of real-world video scenes, and leverage it to enhance spatio-temporal understanding of videos? This issue is particularly important, as many videos originate from real-world audio-visual events, which play a vital role in practical applications, such as human-computer interaction, autonomous driving, and so on, where audio and visual modalities are both crucial for temporal and spatial perception.
@@
In response to this issue, existing datasets such as AVE (Tian et al. 2018), UnAV-100 (Geng et al. 2023), and PU-VALOR (Tang et al. 2025b), incorporate audio-visual information but primarily focus on temporal understanding while overlooking the spatial properties of audible objects. In contrast, spatio-temporal grounding datasets such as VidSTG (Zhang et al. 2020), HC-STVG (Tang et al. 2021), and V-STaR (Cheng et al. 2025) offer both temporal and spatial annotations, but they do not adequately capture the rich audio-visual dynamics of real-world scenes and typically involve only a limited range of object types. Therefore, proposing a fine-grained spatio-temporal dataset that jointly integrates audio and visual information is both valuable and urgent, as it can enhance models’ reasoning capabilities in real-world scenarios.
@@
At the model level, models such as LLaVA-ST (Li et al. 2025a), GroundingGPT (Li et al. 2024) and Grounded-VideoLLM (Wang et al. 2024a) have gradually extended their spatio-temporal modeling capabilities. However, these models require extensive high-quality labeled data and lack sufficient exploration capacity. With the promising performance of Deepseek-R1 (Guo et al. 2025) in rule-based reinforcement learning (RL), some studies have begun exploring the use of GRPO algorithms (Shao et al. 2024) to enhance the reasoning capabilities of MLLMs. Models like VideoChat-R1 (Li et al. 2025b), Video-R1 (Feng et al. 2025) and Omni-R1 (Zhong et al. 2025) have emerged in this direction, yet they offer limited reward designs for spatio-temporal reasoning and lack dedicated tasks tailored for complex audio-visual scenarios.
@@
To address the above limitations, we propose R-AVST, the first video dataset with fine-grained spatio-temporal annotations in complex audio-visual scenarios, as illustrated in Fig. 1. In audio-visual scenarios, both auditory and visual attributes of objects can capture human perception and attention. Motivated by this, we use GPT-4o-mini (Hurst et al. 2024) to extract and analyze audio-visual event captions to label objects’ attributes. This allows models to determine whether an object is audible, visible, or both, and subsequently supports automatic spatial annotation for grounding object locations. To further assess models’ reasoning capabilities and align with human inference needs in complex audio-visual scenes, we define three targeted reasoning tasks, including temporal localization of sounding-visible objects, spatial localization of sounding-visible or silent-visible objects within a given duration, and spatio-temporal localization of sounding-visible objects. Based on these tasks, we automatically generate corresponding question-answer pairs (QAs) to enable large-scale evaluation. In total, R-AVST comprises 5,237 videos with 27,253 objects both sounding-visible and silent-visible and 8,166 QAs, covering over 100 types of audio-visual events, such as human speech, musical performances, and animal sounds.
@@
Building upon R-AVST, we further focus on addressing the challenge of fine-grained spatio-temporal reasoning capabilities of models in complex audio-visual scenarios. To address the lack of advanced reasoning capabilities and the dependence on large-scale high-quality annotated data in models such as LLaVA-ST (Li et al. 2025a), we fine-tune our model using the data-efficient GRPO (Shao et al. 2024) method from DeepSeek-R1 (Guo et al. 2025), given the rule-based characteristics of our tasks. Meanwhile, to overcome the absence of task-specific objectives and reward designs for complex audio-visual spatio-temporal reasoning in models like VideoChat-R1 (Li et al. 2025b), we develop a multi-dimensional reward system tailored to our tasks. This system includes format, object, temporal, and spatial reward, which collectively enable effective policy gradient updates. Our experiments demonstrate that AVST-Zero achieves competitive performance on the three core tasks. It surpasses most Video-LLMs and sets a new perspective in the audio-visual spatio-temporal reasoning tasks.
@@
We collect videos from UnAV-100 (Geng et al. 2023), an audio-visual dataset of untrimmed videos covering diverse domains. Starting from raw YouTube videos and the corresponding event captions, we employ a filtering strategy to select high-quality samples, aiming to ensure comprehensive coverage of complex audio-visual scenarios. As detailed in Fig. 2, the filtering involves three steps to balance video duration, event count, and event coverage ratio. First, videos are grouped into short (0-20s), medium (20-40s), and long (40-60s) durations, excluding extremely short clips and merging overlapping events. Second, we limit videos to at most three audio-visual events, ensuring a balanced distribution of 1, 2, and 3-event videos to facilitate clearer scene-level evaluation. Third, videos with an event-to-total (ET) duration ratio below 0.08 are removed, retaining those where meaningful events occupy a substantial portion. Finally, we curate a dataset of 5,237 high-quality videos, encompassing a wide variety of complex audio-visual scenarios.
@@
Audio-visual spatio-temporal reasoning tasks typically focus on objects, mostly expressed as nouns in captions. We use GPT-4o-mini (Hurst et al. 2024) as an Analyzer LLM to extract noun-based objects from captions, as shown in Fig. 2. To capture the multi-modal nature, tailored prompts guide the model to annotate each object’s auditory and visual attributes in a standardized format. We query the analyzer by emphasizing the definition of audibility in object-sound relations to improve analysis accuracy. For example, in the caption “A group of people are sailing on silver sailboats”, “a group of people” is labeled “visible&audible” while “silver sailboats” is “visible-only”. Finally, we identify 27,253 objects in audio-visual event captions in total, with 50.88% labeled in “visible&audible”.
@@
Based on caption analysis and temporal annotations, we perform spatial annotation on video frames within audio-visual event segments. To reduce the high annotation cost of large-scale videos, we leverage the automatic tool, Grounded-SAM2 (Ravi et al. 2024), for fine-grained, frame-by-frame object annotation. The pipeline extracts frames corresponding to each event and constructs textual prompts based on object information derived from caption analysis, which are then fed into Grounded-SAM2 (Ravi et al. 2024).
@@
As shown in Tab. 1, we compare R-AVST with existing relevant datasets (Gao et al. 2017; Tian et al. 2018; Tang et al. 2025b; Geng et al. 2023, 2025; Zhang et al. 2020; Tang et al. 2021; Yao et al. 2025; Cheng et al. 2025; Zhou et al. 2022, 2025; Liu et al. 2024; Chen et al. 2024a). Existing datasets primarily focus on general temporal or visual grounding scenarios and lack fine-grained modeling of audio-visual scenarios. On the other hand, while datasets such as AVSBench (Zhou et al. 2022, 2025), VPO (Chen et al. 2023), and LU-AVS (Liu et al. 2024) focus on spatial localization of audible objects, they often neglect the temporal aspect. In contrast, R-AVST covers a wider range of events and objects, and is tailored for fine-grained spatio-temporal reasoning in real-world audio-visual scenes, offering a more comprehensive benchmark for complex video reasoning tasks.
@@
To further enhance the spatio-temporal reasoning capabilities of Video-LLMs, we fine-tune the model with GRPO (Shao et al. 2024), along with task-oriented rewards designed to improve performance on the audio-visual spatio-temporal reasoning domain.
@@
GRPO is a reinforcement learning (RL) method and can be used in multimodal large language models (MLLMs) to guide task alignment through reward functions. It differs from PPO in that it reduces reliance on a critic model by directly comparing groups of generated responses. Specifically, for each question $q$, GRPO (Shao et al. 2024) samples a set of outputs $\left\{o_{1},o_{2},\ldots,o_{e}\right\}$ from the old policy $\pi_{\theta_{\text{old }}}$, and then optimizes the policy $\pi_{\theta}$ by maximizing the objective defined as:
@@
We use Word2Vec (Mikolov et al. 2013) to quantify the semantic similarity between the predicted and ground truth object names, which helps mitigate misclassification caused by lexical variations. The similarity $\operatorname{sim}\left(V_{\text{pred}},V_{\text{gt}}\right)$ is defined as:
@@
In experiments, we base Qwen2.5-VL 7B (Bai et al. 2025) and Qwen2.5-Omni 7B (Xu et al. 2025) to fine-tune two model variants on our R-AVST dataset, namely AVST-Zero and AVST-Zero-Omni, respectively. Training is conducted on four NVIDIA RTX A6000 GPUs for a single epoch, with a batch size of 1 on each device. The group generation number is set to 6.
@@
As shown in Tab. 2 and Tab. 3, we compare our models with existing advanced models (Bai et al. 2025; Xu et al. 2025; Cheng et al. 2024; Li et al. 2024; Chen et al. 2024b; Li et al. 2025b; Feng et al. 2025). The evaluation is conducted on the R-AVST test set. In Tab. 2, for the audio-visual spatial reasoning task, AVST-Zero performs similarly to other models in m_vIoU but slightly outperforms them in AP@0.3 with 3.12%. For the audio-visual spatio-temporal reasoning task, AVST-Zero shows significantly better performance, achieving 46.04% m_tIoU and 8.59% m_vIoU. As shown in Tab. 3, AVST-Zero leads in average temporal perception for the audio-visual temporal reasoning task with an m_tIoU of 47.96%. We also observe that AVST-Zero-Omni achieves higher prediction accuracy than AVST-Zero in the object and spatial dimensions, while performing worse in the temporal dimension. This is attributed to the base model’s strong audio–visual joint perception but relatively weak temporal perception capabilities. These results show that our AVST-Zero variants perform competitively across all three tasks, confirming their effectiveness.
@@
As shown in Fig. 5, Qwen2.5-VL (Bai et al. 2025) predicts sparse and inaccurate object locations. VideoChat-R1 (Li et al. 2025b) correctly identifies the girl, but our model yields results closer to the ground truth, with more accurate object recognition and spatial localization.
@@ Acknowledgments
Due to the high cost and inefficiency of manually annotating large-scale video data, we introduce an automated annotation pipeline based on Grounded-SAM2 (Ravi et al. 2024) to perform fine-grained, frame-by-frame object spatio-temporal annotation in complex audio-visual scenarios. In this automated process, we first extract video frames corresponding to the audio-visual events, and then construct text prompts using object names identified during the caption analysis stage. These prompts are used as inputs to Grounded-SAM2 (Ravi et al. 2024). For the configuration, we set the BOX_THRESHOLD to 0.4 and the TEXT_THRESHOLD to 0.3, representing the minimum confidence score for bounding boxes and the minimum similarity score between output text labels and input prompts, respectively. Finally, we obtain bounding box annotations for each frame, which are stored in a structured JSON format.
@@
Following prior works (Yao et al. 2025; Wang et al. 2025; Cheng et al. 2025), we adopt evaluation metrics across different dimensions. At the object level, prediction accuracy is determined by the proportion of cases where the similarity score surpasses a given threshold, as specified in Eq. 2 and Eq. 3. At the temporal level, we evaluate performance using the mean temporal Intersection over Union (m_tIoU) and Recall@1 at tIoU thresholds of {0.3, 0.5, 0.7}. At the spatial level, we evaluate performance using the mean visual IoU (m_vIoU) by frames, and the Average Precision Score (AP)@1 at vIoU thresholds of {0.3, 0.5}.
@@
We employ GPT-4o-mini (Achiam et al. 2023) as our caption analyzer to extract objects and their audio-visual attributes from the audio-visual event captions. As shown in the Box 1, we refine the definition of audible elements and enhance the association between subject and action to guide the model toward more accurate understanding.
@@
In this section, we conduct cross-dataset experiments on additional datasets to further validate the effectiveness of our model. For the temporal reasoning evaluation, we use the AVE (Tian et al. 2018) dataset, which consists of 10-second video clips. For the spatial reasoning evaluation, we adopt the AVSBench-V1 (Zhou et al. 2022) dataset, which contains 5-second videos. As shown in Tab. 6, our model achieves performance comparable to VideoChat-R1 (Li et al. 2025b) in the temporal dimension and attains the best performance in the spatial dimension.
@@
Table 6: Comparison of AVST-Zero on the AVE (Tian et al. 2018) and AVSBench-V1 (Zhou et al. 2022) datasets.

### v2

@@ Introduction
The rapid advancement of multimodal large language models (MLLMs) recently has demonstrated their effectiveness in video understanding by integrating information from various modalities (wang2024comprehensive; zhang2024mm; tang2025video; geng2025longvale). However, this rapid progress raises a question: Do current models and datasets adequately account for the compositional audio-visual nature of real-world video scenes, and leverage it to enhance spatio-temporal understanding of videos? This issue is particularly important, as many videos originate from real-world audio-visual events, which play a vital role in practical applications, such as human-computer interaction, autonomous driving, and so on, where audio and visual modalities are both crucial for temporal and spatial perception.
@@
In response to this issue, existing datasets such as AVE (tian2018audio), UnAV-100 (geng2023dense), and PU-VALOR (tang2025empowering), incorporate audio-visual information but primarily focus on temporal understanding while overlooking the spatial properties of audible objects. In contrast, spatio-temporal grounding datasets such as VidSTG (zhang2020does), HC-STVG (tang2021human), and V-STaR (cheng2025v) offer both temporal and spatial annotations, but they do not adequately capture the rich audio-visual dynamics of real-world scenes and typically involve only a limited range of object types. Therefore, proposing a fine-grained spatio-temporal dataset that jointly integrates audio and visual information is both valuable and urgent, as it can enhance models’ reasoning capabilities in real-world scenarios.
@@
At the model level, models such as LLaVA-ST (li2025llava), GroundingGPT (li2024groundinggpt) and Grounded-VideoLLM (wang2024grounded) have gradually extended their spatio-temporal modeling capabilities. However, these models require extensive high-quality labeled data and lack sufficient exploration capacity. With the promising performance of Deepseek-R1 (guo2025deepseek) in rule-based reinforcement learning (RL), some studies have begun exploring the use of GRPO algorithms (shao2024deepseekmath) to enhance the reasoning capabilities of MLLMs. Models like VideoChat-R1 (li2025videochat), Video-R1 (feng2025video) and Omni-R1 (zhong2025omni) have emerged in this direction, yet they offer limited reward designs for spatio-temporal reasoning and lack dedicated tasks tailored for complex audio-visual scenarios.
@@
To address the above limitations, we propose R-AVST, the first video dataset with fine-grained spatio-temporal annotations in complex audio-visual scenarios, as illustrated in Fig. 1. In audio-visual scenarios, both auditory and visual attributes of objects can capture human perception and attention. Motivated by this, we use GPT-4o-mini (hurst2024gpt) to extract and analyze audio-visual event captions to label objects’ attributes. This allows models to determine whether an object is audible, visible, or both, and subsequently supports automatic spatial annotation for grounding object locations. To further assess models’ reasoning capabilities and align with human inference needs in complex audio-visual scenes, we define three targeted reasoning tasks, including temporal localization of sounding-visible objects, spatial localization of sounding-visible or silent-visible objects within a given duration, and spatio-temporal localization of sounding-visible objects. Based on these tasks, we automatically generate corresponding question-answer pairs (QAs) to enable large-scale evaluation. In total, R-AVST comprises 5,237 videos with 27,253 objects both sounding-visible and silent-visible and 8,166 QAs, covering over 100 types of audio-visual events, such as human speech, musical performances, and animal sounds.
@@
Building upon R-AVST, we further focus on addressing the challenge of fine-grained spatio-temporal reasoning capabilities of models in complex audio-visual scenarios. To address the lack of advanced reasoning capabilities and the dependence on large-scale high-quality annotated data in models such as LLaVA-ST (li2025llava), we fine-tune our model using the data-efficient GRPO (shao2024deepseekmath) method from DeepSeek-R1 (guo2025deepseek), given the rule-based characteristics of our tasks. Meanwhile, to overcome the absence of task-specific objectives and reward designs for complex audio-visual spatio-temporal reasoning in models like VideoChat-R1 (li2025videochat), we develop a multi-dimensional reward system tailored to our tasks. This system includes format, object, temporal, and spatial reward, which collectively enable effective policy gradient updates. Our experiments demonstrate that AVST-Zero achieves competitive performance on the three core tasks. It surpasses most Video-LLMs and sets a new perspective in the audio-visual spatio-temporal reasoning tasks.
@@
Spatio-Temporal Understanding in Video-LLMs
Spatio-temporal understanding is crucial for extracting key information from videos (goodge2025spatio; yuan2024video; zou2023beyond; geng2024uniav). With the rise of MLLMs, general-purpose Video-LLMs like InternVL-2.5 (chen2024expanding), Qwen2.5-VL (bai2025qwen2), VideoLLaMA3 (zhang2025videollama), and GroundingGPT (li2024groundinggpt) have made notable strides, along with specialized models such as LLaVA-ST (li2025llava), VideoMolmo (ahmad2025videomolmo), and Meerkat (chowdhury2024meerkat) that focus on spatio-temporal reasoning. These models benefit from improvements like stronger encoders and audio integration, yet still struggle with complex audio-visual spatio-temporal tasks. On the dataset side, benchmarks like HC-STVG (tang2021human) and VidSTG (zhang2020does) provide spatio-temporal grounding annotations, but feature limited object diversity. While BOSTVG (yao2025omnistvg) introduces a multi-object setting but uses short videos. V-STaR (cheng2025v) targets long-video scenarios, yet lacks realistic audio-visual content. AV-UIE (du2025crab) centers on audio-visual understanding but mainly targets image-audio pairs, overlooking the construction of video-audio data.
Reinforcement Learning Enhancement in LLMs
Reinforcement Learning (RL) effectively enhances LLM reasoning with limited supervision. GPT-4 (achiam2023gpt) uses PPO with reward and value functions, while ChatGLM3-DPO (glm2024chatglm) applies DPO through pairwise comparisons. DeepSeek-R1 (guo2025deepseek) introduces GRPO (shao2024deepseekmath), using rewards within groups to improve reasoning without value functions. GRPO has since been adopted in Video-LLMs as a method after supervised fine-tuning (SFT). VideoChat-R1 (li2025videochat) improves spatio-temporal perception, Video-R1 (feng2025video) enhances temporal modeling, and R1-Omni (zhao2025r1) incorporates audio for emotion understanding. Recent work (guo2025deepseek) explores fully RL-based training, as in Omni-R1 (zhong2025omni), which proposes an end-to-end GRPO framework. However, RL models specifically targeting spatio-temporal reasoning in audio-visual scenarios remain underexplored.
@@
We collect videos from UnAV-100 (geng2023dense), an audio-visual dataset of untrimmed videos covering diverse domains. Starting from raw YouTube videos and the corresponding event captions, we employ a filtering strategy to select high-quality samples, aiming to ensure comprehensive coverage of complex audio-visual scenarios. As detailed in Fig. 2, the filtering involves three steps to balance video duration, event count, and event coverage ratio. First, videos are grouped into short (0-20s), medium (20-40s), and long (40-60s) durations, excluding extremely short clips and merging overlapping events. Second, we limit videos to at most three audio-visual events, ensuring a balanced distribution of 1, 2, and 3-event videos to facilitate clearer scene-level evaluation. Third, videos with an event-to-total (ET) duration ratio below 0.08 are removed, retaining those where meaningful events occupy a substantial portion. Finally, we curate a dataset of 5,237 high-quality videos, encompassing a wide variety of complex audio-visual scenarios.
@@
Audio-visual spatio-temporal reasoning tasks typically focus on objects, mostly expressed as nouns in captions. We use GPT-4o-mini (hurst2024gpt) as an Analyzer LLM to extract noun-based objects from captions, as shown in Fig. 2. To capture the multi-modal nature, tailored prompts guide the model to annotate each object’s auditory and visual attributes in a standardized format. We query the analyzer by emphasizing the definition of audibility in object-sound relations to improve analysis accuracy. For example, in the caption “A group of people are sailing on silver sailboats”, “a group of people” is labeled “visible&audible” while “silver sailboats” is “visible-only”. Finally, we identify 27,253 objects in audio-visual event captions in total, with 50.88% labeled in “visible&audible”.
@@
Based on caption analysis and temporal annotations, we perform spatial annotation on video frames within audio-visual event segments. To reduce the high annotation cost of large-scale videos, we leverage the automatic tool, Grounded-SAM2 (ravi2024sam), for fine-grained, frame-by-frame object annotation. The pipeline extracts frames corresponding to each event and constructs textual prompts based on object information derived from caption analysis, which are then fed into Grounded-SAM2 (ravi2024sam).
@@
As shown in Tab. 1, we compare R-AVST with existing relevant datasets (gao2017tall; tian2018audio; tang2025empowering; geng2023dense; geng2025longvale; zhang2020does; tang2021human; yao2025omnistvg; cheng2025v; zhou2022audio; zhou2025audio; liu2024benchmarking; chen2024unraveling). Existing datasets primarily focus on general temporal or visual grounding scenarios and lack fine-grained modeling of audio-visual scenarios. On the other hand, while datasets such as AVSBench (zhou2022audio; zhou2025audio), VPO (chen2023closer), and LU-AVS (liu2024benchmarking) focus on spatial localization of audible objects, they often neglect the temporal aspect. In contrast, R-AVST covers a wider range of events and objects, and is tailored for fine-grained spatio-temporal reasoning in real-world audio-visual scenes, offering a more comprehensive benchmark for complex video reasoning tasks.
@@
To further enhance the spatio-temporal reasoning capabilities of Video-LLMs, we fine-tune the model with GRPO (shao2024deepseekmath), along with task-oriented rewards designed to improve performance on the audio-visual spatio-temporal reasoning domain.
@@
GRPO is a reinforcement learning (RL) method and can be used in multimodal large language models (MLLMs) to guide task alignment through reward functions. It differs from PPO in that it reduces reliance on a critic model by directly comparing groups of generated responses. Specifically, for each question $q$, GRPO (shao2024deepseekmath) samples a set of outputs $\left\{o_{1},o_{2},\ldots,o_{e}\right\}$ from the old policy $\pi_{\theta_{\text{old }}}$, and then optimizes the policy $\pi_{\theta}$ by maximizing the objective defined as:
@@
We use Word2Vec (mikolov2013efficient) to quantify the semantic similarity between the predicted and ground truth object names, which helps mitigate misclassification caused by lexical variations. The similarity $\operatorname{sim}\left(V_{\text{pred}},V_{\text{gt}}\right)$ is defined as:
@@
In experiments, we base Qwen2.5-VL 7B (bai2025qwen2) and Qwen2.5-Omni 7B (xu2025qwen2) to fine-tune two model variants on our R-AVST dataset, namely AVST-Zero and AVST-Zero-Omni, respectively. Training is conducted on four NVIDIA RTX A6000 GPUs for a single epoch, with a batch size of 1 on each device. The group generation number is set to 6.
@@
As shown in Tab. 2 and Tab. 3, we compare our models with existing advanced models (bai2025qwen2; xu2025qwen2; cheng2024videollama; li2024groundinggpt; chen2024expanding; li2025videochat; feng2025video). The evaluation is conducted on the R-AVST test set. In Tab. 2, for the audio-visual spatial reasoning task, AVST-Zero performs similarly to other models in m_vIoU but slightly outperforms them in AP@0.3 with 3.12%. For the audio-visual spatio-temporal reasoning task, AVST-Zero shows significantly better performance, achieving 46.04% m_tIoU and 8.59% m_vIoU. As shown in Tab. 3, AVST-Zero leads in average temporal perception for the audio-visual temporal reasoning task with an m_tIoU of 47.96%. We also observe that AVST-Zero-Omni achieves higher prediction accuracy than AVST-Zero in the object and spatial dimensions, while performing worse in the temporal dimension. This is attributed to the base model’s strong audio–visual joint perception but relatively weak temporal perception capabilities. These results show that our AVST-Zero variants perform competitively across all three tasks, confirming their effectiveness.
@@
As shown in Fig. 5, Qwen2.5-VL (bai2025qwen2) predicts sparse and inaccurate object locations. VideoChat-R1 (li2025videochat) correctly identifies the girl, but our model yields results closer to the ground truth, with more accurate object recognition and spatial localization.
@@ Acknowledgments
Due to the high cost and inefficiency of manually annotating large-scale video data, we introduce an automated annotation pipeline based on Grounded-SAM2 (ravi2024sam) to perform fine-grained, frame-by-frame object spatio-temporal annotation in complex audio-visual scenarios. In this automated process, we first extract video frames corresponding to the audio-visual events, and then construct text prompts using object names identified during the caption analysis stage. These prompts are used as inputs to Grounded-SAM2 (ravi2024sam). For the configuration, we set the BOX_THRESHOLD to 0.4 and the TEXT_THRESHOLD to 0.3, representing the minimum confidence score for bounding boxes and the minimum similarity score between output text labels and input prompts, respectively. Finally, we obtain bounding box annotations for each frame, which are stored in a structured JSON format.
@@
Following prior works (yao2025omnistvg; wang2025spacevllm; cheng2025v), we adopt evaluation metrics across different dimensions. At the object level, prediction accuracy is determined by the proportion of cases where the similarity score surpasses a given threshold, as specified in Eq. 2 and Eq. 3. At the temporal level, we evaluate performance using the mean temporal Intersection over Union (m_tIoU) and Recall@1 at tIoU thresholds of {0.3, 0.5, 0.7}. At the spatial level, we evaluate performance using the mean visual IoU (m_vIoU) by frames, and the Average Precision Score (AP)@1 at vIoU thresholds of {0.3, 0.5}.
@@
We employ GPT-4o-mini (achiam2023gpt) as our caption analyzer to extract objects and their audio-visual attributes from the audio-visual event captions. As shown in the Box 1, we refine the definition of audible elements and enhance the association between subject and action to guide the model toward more accurate understanding.
@@
In this section, we conduct cross-dataset experiments on additional datasets to further validate the effectiveness of our model. For the temporal reasoning evaluation, we use the AVE (tian2018audio) dataset, which consists of 10-second video clips. For the spatial reasoning evaluation, we adopt the AVSBench-V1 (zhou2022audio) dataset, which contains 5-second videos. As shown in Tab. 6, our model achieves performance comparable to VideoChat-R1 (li2025videochat) in the temporal dimension and attains the best performance in the spatial dimension.
@@
Table 6: Comparison of AVST-Zero on the AVE (tian2018audio) and AVSBench-V1 (zhou2022audio) datasets.